# One-dimensional long-range Ising model: two (almost) equivalent approximations

### v1

@@ I Introduction
where $\sigma>0$ ensures a well-defined thermodynamics [7]. Unlike its short-range counterpart in $d=1$, the 1D LRI model exhibits a finite-temperature phase transition for certain values of $\sigma$. This was first rigorously established by Dyson in Ref. [8] for all $\sigma\in(0,1)$, through the introduction of Dyson’s hierarchical model (DHM) as a tractable lower bound. For $\sigma>1$ there is no transition [9]. The marginal case $\sigma=1$, corresponding to interactions decaying as $1/r^{2}$, remained unresolved in Dyson’s original analysis. It was later shown by Fröhlich and Spencer [10] that this case features a Berezinskii–Kosterlitz–Thouless (BKT)-like transition, characterized by an essential singularity in the free energy. Notably, however, this transition includes a discontinuity in the magnetization – anticipated by Thouless in Ref. [11], see also [12] – which distinguishes it from the standard BKT transition in the two-dimensional XY model. The strong interest in the Ising model with $1/r^{2}$ interactions was the result of its equivalence to the Kondo problem, established by Anderson, Yuval and Hamann in a series of papers [13, 14, 15], which also contain an early version of the renormalization group (RG) later developed by Wilson [16]. Their methods were extended by Cardy [17] to encompass general discrete spin chains with $1/r^{2}$ interactions, including Potts and Ashkin–Teller models.
@@
In this work, we adopt the RG framework, which has played a central role in the study of long-range critical systems [27, 28]. Our goal is twofold. First, to compute the critical exponent $\nu$ within a functional renormalization group (FRG) approach [29, 30, 31, 32] – notice, indeed, that one expects that the critical exponent $\eta$ has a trivial dependence on $\sigma$ [28] and therefore the critical exponents are determined once the non-trivial dependence of $\nu$ has been worked out. Second, to explore the connection between the FRG analysis of the 1D LRI model and the real-space RG treatment of Dyson’s hierarchical model, which is a well-developed tool [33, 34]. Although the DHM is not translationally invariant and is often regarded as a toy model, we argue that it admits an effective description that closely mirrors the local potential approximation (LPA) of the FRG.
@@ II Reminder on critical properties of long-range systems
Focusing on the regime $\sigma>0$, where interactions decay more rapidly than $1/r^{d}$, we distinguish between three scenarios. In the interval $\sigma\in(0,\sigma_{\rm mf})$ the critical behavior is essentially mean-field and governed by the Gaussian fixed point, very similar to how mean-field theory provides the correct exponents above the upper critical dimension in short-range systems. Distinctive long-range features emerge for $\sigma\in(\sigma_{\rm mf},\sigma_{*})$, where the critical exponents depend continuously on the range parameter $\sigma$. In this intermediate regime – the focus of this work – the critical fixed points are interacting ones, meaning that the role of fluctuations becomes crucial. Finally, for $\sigma\in(\sigma_{*},\infty)$ the critical behavior crosses over to the short-range universality class, as the long-range interaction becomes irrelevant. While $\sigma_{\rm mf}=d/2$ for classical LR $O(n)$ models [1, 7], the exact value of $\sigma_{*}$ has long remained controversial in d>1d>1[35]. Today, the most widely accepted scenario, proposed by Sak [28], is that $\sigma_{*}=2-\eta_{\rm SR}$, where $\eta_{\rm SR}$ is the anomalous dimension of the model with short-range interactions, while for the long-range model $\eta(\sigma)=2-\sigma$[22, 36, 37]. Critical properties of $d$-dimensional long-range systems have been studied by a variety of methods, including Monte Carlo [38, 39, 22, 40], RG approaches [27, 28, 23, 24, 36] and conformal bootstrap [37, 41, 42, 43]; see further references in [7, 35, 1]. Peculiar behavior may arise in two-dimensional $O(2)$ symmetric models [44, 45, 46, 47, 48, 49], which is yet the subject of active research.
@@ III.1 Functional renormalization group in the LPA
proposed by Wetterich in [29], where $\int_{q}=\int dq/(2\pi)$ and $t=\log(k/\Lambda)$. Denoting the second functional derivative (Hessian) of the action by $\Gamma_{k}^{(2)}$, the field-dependent propagator $G_{k}=G_{k}[\phi]$ reads
@@ III.2 RG flow and fixed points
that generalizes the Litim cutoff used for short-range systems [56, 57, 58]. This choice is discussed more carefully in Sec. IV.3. Here we comment that in general the Litim regulator within LPA gives good results for critical exponents [32], and, most importantly for our purposes, is the one that highlights the connection with the DHM studied in the next section. Then, Eq. (6) becomes
@@
Setting the left-hand side of (9) to zero yields the fixed-point equation. This is now an ordinary differential equation whose solutions $\tilde{V}_{*}(\varphi)$ can be studied by means of a shooting method known as spikeplot [59, 60, 36]. Both the Gaussian and the interacting (Wilson-Fisher) fixed points are found without the need of any Taylor expansion of the effective potential.
@@ III.3 Linearization around the fixed point
Figure 1: Inverse of the critical exponent $\nu$ of the 1D LRI model in the region $1/2<\sigma<1$. The red dots are the results coming from a non-perturbative FRG calculation at the LPA level, described in Sec. III. The blue crosses represent the estimates obtained from the DHM via a real-space RG, see Sec. IV. The green triangles at $\sigma=0.5$, $\sigma\approx 0.65$ and $\sigma=0.875$ are obtained via the effective dimension approach, Eq. (47). The dotted line is the two-loop $\epsilon$ expansion (44) of [27]. The black dashed and olive dash-dotted lines are the results coming from expansions around $\sigma=1$, respectively from [23] and [25]. More details about these are provided in Sec. V. A detailed comparison of LPA and DHM results is presented in Fig. 5. Finally, Monte Carlo points are taken from Refs. [21] and  [20].
@@ IV Hierarchical model
Before proceeding to discuss the renormalization of the DHM, we remark that the model displays a phase transition for all values of $\alpha\in(1,2)$ – or $\sigma\in(0,1)$ – like the 1D LRI model [8]. The case $\sigma=1$ is more delicate, because the 1D LRI model has a phase transition with BKT features [11, 10], while the DHM does not [8]. Nonetheless, a slight modification of the decay of the interaction in Eq. (14) – namely replacing the factor $2^{-p(1+\sigma)}=2^{-2p}$ by $2^{-2p}\log(p)$does yield a BKT-like transition and the Thouless effect [12].
@@ IV.3 Exponent $\nu$ and comparison with the LPA
with $l=-k+k^{\prime}+k^{\prime\prime}$. When the temperature is tuned to its critical value $\beta=\beta_{c}$, entering the initial condition $\tilde{P}_{0}$ and thus the coefficients $\{B_{k}^{(p=0)}\}$, one is able to reach a fixed point characterized by $\{B_{k}^{*}\}$ and $\mu^{*}$, which is the obvious notation for (39) with $B_{k}^{(p)}\to B_{k}^{*}$. The fixed point can be found with very high accuracy and with a truncated series (37), which typically need not include much more than $10$ terms [33].
@@
The critical exponent $\nu$ is obtained from the largest eigenvalue $y_{1}$ of the stability matrix,
$\nu^{-1}=\log y_{1}/\log 2,$ (43)
@@
Figure 5: In the upper panel we show the thermal eigenvalue $y=\nu^{-1}$ obtained as solution of (12) in the LPA and (43) for the DHM. The points overlap almost exactly, as in Fig. 1. The relative difference $\Delta y=y_{\rm LPA}-y_{\rm DHM}$, where $y_{\rm mean}=(y_{\rm LPA}+y_{\rm DHM})/2$, is shown in the lower panel. The error is peaked in the region of large $\sigma$, but it stays below $10^{-3}$ throughout the interval $1/2<\sigma<1$.
@@ VI Conclusions
Looking forward, several interesting directions emerge. First, it would be desirable to improve upon the LPA by incorporating the effects of short-range analytic terms and testing the robustness of the observed equivalence beyond leading-order truncations. Additionally, adapting the FRG framework to study the effective DHM field theory – with its hierarchical Laplacian dispersion $\omega(q)$ – may allow for a direct and controlled comparison with numerical RG flows. From a broader perspective, the methods and analogies presented here may inform the treatment of more complex systems exhibiting long-range interactions, including quantum spin chains and out-of-equilibrium models.

### v2

@@ I Introduction
where $\sigma>0$ ensures a well-defined thermodynamics [7]. Unlike its short-range counterpart in $d=1$, the 1D LRI model exhibits a finite-temperature phase transition for certain values of $\sigma$. In fact, the long-range nature of the couplings suppresses the strong fluctuations typical of low-dimensional systems, allowing for spontaneous symmetry breaking. This was first rigorously established by Dyson in Ref. [8] for all $\sigma\in(0,1)$, through the introduction of Dyson’s hierarchical model (DHM) as a tractable lower bound. For $\sigma>1$ there is no transition [9]. The marginal case $\sigma=1$, corresponding to interactions decaying as $1/r^{2}$, remained unresolved in Dyson’s original analysis. It was later shown by Fröhlich and Spencer [10] that this case features a Berezinskii–Kosterlitz–Thouless (BKT)-like transition, characterized by an essential singularity in the free energy. Notably, however, this transition includes a discontinuity in the magnetization – anticipated by Thouless in Ref. [11], see also [12] – which distinguishes it from the standard BKT transition in the two-dimensional XY model. The strong interest in the Ising model with $1/r^{2}$ interactions was the result of its equivalence to the Kondo problem, established by Anderson, Yuval and Hamann in a series of papers [13, 14, 15], which also contain an early version of the renormalization group (RG) later developed by Wilson [16]. Their methods were extended by Cardy [17] to encompass general discrete spin chains with $1/r^{2}$ interactions, including Potts and Ashkin–Teller models.
@@
In this work, we adopt the RG framework, which has played a central role in the study of long-range critical systems [27, 28]. Our goal is twofold. First, to compute the critical exponent $\nu$ within a functional renormalization group (FRG) approach [29, 30, 31, 32] – notice, indeed, that one expects that the critical exponent $\eta$ has a trivial dependence on $\sigma$ [28] and therefore the main critical exponents are determined once the non-trivial dependence of $\nu$ has been worked out. Second, to explore the connection between the FRG analysis of the 1D LRI model and the real-space RG treatment of Dyson’s hierarchical model, which is a well-developed tool [33, 34], even from a mathematically rigorous point of view [35, 36, 37]. Although the DHM is not translationally invariant and is often regarded as a toy model, we argue that it admits an effective description that closely mirrors the local potential approximation (LPA) of the FRG.
@@ II Reminder on critical properties of long-range systems
Focusing on the regime $\sigma>0$, where interactions decay more rapidly than $1/r^{d}$, we distinguish between three scenarios. In the interval $\sigma\in(0,\sigma_{\rm mf})$ the critical behavior is essentially mean-field and governed by the Gaussian fixed point, very similar to how mean-field theory provides the correct exponents above the upper critical dimension in short-range systems. Distinctive long-range features emerge for $\sigma\in(\sigma_{\rm mf},\sigma_{*})$, where the critical exponents depend continuously on the range parameter $\sigma$. In this intermediate regime – the focus of this work – the critical fixed points are interacting ones, meaning that the role of fluctuations becomes crucial. Finally, for $\sigma\in(\sigma_{*},\infty)$ the critical behavior crosses over to the short-range universality class, as the long-range interaction becomes irrelevant. While $\sigma_{\rm mf}=d/2$ for classical LR $O(n)$ models [1, 7], the exact value of $\sigma_{*}$ has long remained controversial in d>1d>1[38]. Today, the most widely accepted scenario, proposed by Sak [28], is that $\sigma_{*}=2-\eta_{\rm SR}$, where $\eta_{\rm SR}$ is the anomalous dimension of the model with short-range interactions, while for the long-range model $\eta(\sigma)=2-\sigma$[22, 39, 40]. Critical properties of $d$-dimensional long-range systems have been studied by a variety of methods, including Monte Carlo [41, 42, 22, 43], RG approaches [27, 28, 23, 24, 39] and conformal bootstrap [40, 44, 45, 46]; see further references in [7, 38, 1]. Peculiar behavior may arise in two-dimensional $O(2)$ symmetric models [47, 48, 49, 50, 51], which is yet the subject of active research.
@@ III.1 Functional renormalization group in the LPA
proposed by Wetterich and Morris [29, 54], where $\int_{q}=\int dq/(2\pi)$ and $t=\log(k/\Lambda)$. Denoting the second functional derivative (Hessian) of the action by $\Gamma_{k}^{(2)}$, the field-dependent propagator $G_{k}=G_{k}[\phi]$ reads
@@ III.2 RG flow and fixed points
generalizing the standard Litim cutoff used for short-range systems. We employ this regulator because it satisfies an optimization criterion for the LPA [59, 60, 61], thereby minimizing systematic errors in the computation of critical exponents [32]. For example, the critical exponents found with the Litim regulator are more accurate [62] than those obtained with a power-law cutoff [63, 64]. Most importantly for our purposes, this optimized choice is structurally closest to the recursion relations of the DHM, allowing for the direct comparison performed in Sec. IV.3. The investigation of different regulator choices and their impact on the specific numerical values of critical exponents is deferred to future studies. Focusing here on the optimized Litim cutoff, the flow equation Eq. (6) reads
@@
Setting the left-hand side of (9) to zero yields the fixed-point equation. This is now an ordinary differential equation whose solutions $\tilde{V}_{*}(\varphi)$ can be studied by means of a shooting method known as spikeplot [65, 66, 39] or via pseudospectral approaches [67, 68]. Both the Gaussian and the interacting (Wilson-Fisher) fixed points are found without the need of any Taylor expansion of the effective potential.
@@ III.3 Linearization around the fixed point
Figure 1: Inverse of the critical exponent $\nu$ of the 1D LRI model in the region $1/2<\sigma<1$. The red dots are the results coming from a non-perturbative FRG calculation at the LPA level, described in Sec. III. The blue crosses represent the estimates obtained from the DHM via a real-space RG, see Sec. IV. The green triangles at $\sigma=0.5$, $\sigma\approx 0.65$ and $\sigma=0.875$ are obtained via the effective dimension approach, Eq. (47). The dotted line is the two-loop $\epsilon$ expansion (44) of [27], while the pink stars are the three-loop Padé-Borel resummed exponents of [69]. The black dashed and olive dash-dotted lines are the results coming from expansions around $\sigma=1$, respectively from [23] and [25]. More details about these are provided in Sec. V. A detailed comparison of LPA and DHM results is presented in Fig. 5. Finally, Monte Carlo points are taken from Refs. [21] and  [20].
@@ IV Hierarchical model
We remark in passing that the field of p-adic numbers, which can be represented as infinite paths in a rooted tree (in our case a 2-adic tree), plays an important role in the rigorous RG-based formulation of quantum field theory, see e.g. [72, 73].
@@
The self-similar nature of hierarchical tree structures can be used to derive exact RG equations in several different context, from critical phenomena [33] to polymers [74], information transitions [75] and approximated ones for quantum phase transitions [76]. Before proceeding to discuss the renormalization of the classical DHM, we remark that the model displays a phase transition for all values of $\alpha\in(1,2)$ – or $\sigma\in(0,1)$ – like the 1D LRI model [8]. The case $\sigma=1$ is more delicate, because the 1D LRI model has a phase transition with BKT features [11, 10], while the DHM does not [8]. Nonetheless, a slight modification of the decay of the interaction in Eq. (14) – namely replacing the factor $2^{-p(1+\sigma)}=2^{-2p}$ by $2^{-2p}\log(p)$does yield a BKT-like transition and the Thouless effect [12]. We leave a more careful study of this variation of the DHM to future investigations.
@@ IV.3 Exponent $\nu$ and comparison with the LPA
with $l=-k+k^{\prime}+k^{\prime\prime}$. When the temperature is tuned to its critical value $\beta=\beta_{c}$, entering the initial condition $\tilde{P}_{0}$ and thus the coefficients $\{B_{k}^{(p=0)}\}$, one is able to reach a fixed point characterized by $\{B_{k}^{*}\}$ and $\mu^{*}$, which is the obvious notation for (39) with $B_{k}^{(p)}\to B_{k}^{*}$. The fixed point can be found with very high accuracy and with a truncated series (37), which typically need not include much more than $10$ terms [33]. The rigorous construction of this interacting fixed point, along with further aspects of the RG flow of Dyson-type hierarchical models, is explored in [35, 36, 37].
@@
The critical exponent $\nu$ is obtained from the largest eigenvalue $\lambda_{1}$ of the stability matrix,
$\nu^{-1}=\log\lambda_{1}/\log 2,$ (43)
@@
Figure 5: In the upper panel we show the thermal eigenvalue $y=\nu^{-1}$ obtained as solution of (12) in the LPA and (43) for the DHM. The points overlap almost exactly, as in Fig. 1. Their relative difference, where $\Delta y=y_{\rm LPA}-y_{\rm DHM}$ and $y_{\rm mean}=(y_{\rm LPA}+y_{\rm DHM})/2$, is shown in the lower panel. The error is peaked in the region of large $\sigma$, but it stays below $10^{-3}$ throughout the interval $1/2<\sigma<1$.
@@
For completeness, we also report in Fig. 6 a comprehensive picture of the first few RG eigenvalues, where $y_{1}=\nu^{-1}$ and $\omega=-y_{2}$ is the so-called ‘correction-to-scaling’ exponent. In the region $0<\sigma\leq 1/2$ the Gaussian fixed point dictates the universal properties of the system and the eigenvalues are given by $y_{n}=n(\sigma-1)+1$ [33]. As we cross $\sigma=1/2$ towards larger values, as in Fig. 5, the Gaussian fixed point exchanges stability with the interacting one, whose RG eigenvalues are computed via FRG (filled circles) or the method of [33] (crosses). Not only the thermal eigenvalue $y_{1}$, but also the eigenvalues $y_{n>1}$ corresponding to irrelevant perturbations exhibit the LPA-DHM correspondence described previously. The discrepancy between the two methods is quantitatively comparable to that for $y_{1}$.
Figure 6: First three RG eigenvalues $y_{1},y_{2},y_{3}$ computed by means of the FRG of Sec. III (filled circles) or the RG for the DHM of Sec. IV.3 (crosses). We find close agreement between the latter methods around the interacting fixed-point, while the eigenvalues are exact (solid lines) for the Gaussian region $0<\sigma<1/2$.
@@ V Comparison with other methods
A more thorough weak-coupling expansion was carried out for long-range models in [81, 69]. This work provides three-loop estimates of the critical exponents in the $\epsilon$-expansion. In order to obtain accurate results, one additionally needs to perform a Padé-Borel resummation of the three-loop series. The results are plotted in Fig. 1 as pink stars. At $\sigma=0.6$, these higher-order corrections realign the perturbative results with the LPA, while at $\sigma=0.7$ and $0.8$, the exponent $\nu^{-1}$ becomes compatible with [21] and larger than the LPA one.
@@ VI Conclusions
Looking forward, several interesting directions emerge. First, it would be desirable to improve upon the LPA by incorporating the effects of short-range analytic terms and testing the robustness of the observed equivalence beyond leading-order truncations. Additionally, adapting the FRG framework to study the effective DHM field theory – with its hierarchical Laplacian dispersion $\omega(q)$ – may allow for a direct and controlled comparison with numerical RG flows. From a broader perspective, the methods and analogies presented here may inform the treatment of more complex systems exhibiting long-range interactions, including quantum spin chains and out-of-equilibrium models [6, 85].
@@
Note added. During the completion of this work, we became aware of a related preprint [86], analyzing the 1D LRI model via real-space RG, and reporting finite-temperature phases for the ferromagnetic and spin-glass regimes but no ordered phase in the antiferromagnetic case.